259L oil of compressibility k=20×10^-6 atm^-1is subjected to a hydraulic press. Find the decrease in the value of the oil when it is subjected to pressure 1.6×10^7

1
Expert's answer
2022-01-28T08:14:17-0500

Given:

"V_0=259\:\rm L"

"k=20×10^{-6}\:\rm atm^{-1}"

"\Delta P=1.6\times 10^7\:\rm Pa=157.9\: atm"


The compressibility is given by

"k=-\frac{1}{V_0}\frac{\Delta V}{\Delta P}"

The final volume

"V=V_0+\Delta V=V_0(1-k\Delta P)"

"V=\rm 259\: L(1-20*10^{-6}*159.7)=258\: L"